Three identical bulbs are connected to a cell in series. Which bulb has the largest current through it?


1
Expert's answer
2021-06-22T10:10:25-0400

Gives

Three bulb are series combination


Cell voltage =V

Current="I"


Three resistance

"R_1=R"

"R_2=R"

"P_1=I^2R"

"R_3=R"

"P_2=I^2R"

"P_3=I^2R"


All three resistance are equal and current flow same in series combination then Power

Pnet=P1+P2+P3

"P_{net}=3P"


"P=\frac{P_{net}}{3}"


"P_1=P_2=P_3=P=I^2R"



All three resistance are same current flow

And all three resistance are same power